The cost to rent a car for a certain number of hours is shown in the table. The total cost includes a fixed fee and an hourly rate.
Number of hours Cost
3 $89.25
6 $118.50
14 $196.50
A. Write an equation to model the data. Use y to represent the total cost and x to represent the number of hours the car is rented.
B. What is the fixed fee charged to rent the car?
c. What is the hourly rate charged to rent the car?

Found 2 solutions by stanbon, ewatrrr:
Answer by stanbon(75887)

The cost to rent a car for a certain number of hours is shown in the table. The total cost includes a fixed fee and an hourly rate.
Number of hours Cost
3 $89.25
6 $118.50
14 $196.50
A. Write an equation to model the data. Use y to represent the total cost and x to represent the number of hours the car is rented.
Equation:: y = fixed + h*x
------
Equations based on data:
f + 3h = 89.25
f + 6h = 118.50
f +14h = 196.50
-------------------------
Using the top two equations:
3h = 118.50-89.25
3h = 29.25
h = $9.25
B. What is the fixed fee charged to rent the car?
f + 3h = 89.25
f = 89.25-29.25 = $60.00
--------------------------------
c. What is the hourly rate charged to rent the car?
h = $9.25
----------------------
Cheers,
Stan H.
----------------------

Answer by ewatrrr(24785)

 
Hi,
using:
hrs Cost
3 $89.25
6 $118.50
3h + f = $89.25 |multiplying thru by -2 to eliminate the h variable
6h + f = $118.50
-6h -2f = -$178.50
6h + f = $118.50
f = $60
3h + $60 = $89.25
3h = $29.25
h = $9.75 per hour
y = $9.75x + $60
CHECKING our answer***
$9.75*14 + $60 = $196.50
Wish You the Best in your Studies.